Write the equation of the line that passes through (2,7) and is parallel to the line 2x+3y=18
Answer by checkley75(3666)

2X+3Y=18
3Y=-2X+18
Y=-2X/3+18/3
Y=-2X/3+6 NOW WE HAVE A SLOPE OF -2.
NOW SUBSTITUTE (2,7) FOR X & Y
7=-2/3*2+b
7=-4/3+b
b=7+4/3
b=25/3 Y INTERCEPT
Y=-2X/3+25/3 IS THIS LINES EQUATION.
HERE IS THE GRAPH OF THESE 2 LINES
(graph 300x200 pixels, x from -6 to 5, y from -10 to 10, of TWO functions y = -2x/3 +6 and y = -2x/3 +25/3).